2. Ils pointsy) Explain what liner, ua mean in the term linear quadratic regulator(LQR) question sheet
Answer #1

Hello,
          Please find the answer attached as under. Please give a thumbs up rating if you find the answer useful! Have a rocking day ahead!

The term "linear" means that the control action is linear i.e., it obeys the laws of Additivity and Homogeneity:

  • Additivity: f(x + y) = f(x) + f(y).
  • Homogeneity of degree 1: fx) = αf(x) for all α.

The term "quadratic" is added to imply a quadratic cost function, associated with the control action. A quadratic cost function is taken and the integral of the same is minimised to get the optimal control law. A quadratic cost function has the advantage of being greater than or equal to zero for all values of the control variable.

The term "regulator" simply means "controller". Since this is a feedback control law, it regulates the value of a control variable.

*****************************************************************
PS: Please do not forget the thumbs up!